OpenStudy (hunter68):

Will fan and medal please :) Roller Coaster Crew Ray and Kelsey have summer internships at an engineering firm. As part of their internship, they get to assist in the planning of a brand new roller coaster. For this assignment, you help Ray and Kelsey as they tackle the math behind some simple curves in the coaster's track.

OpenStudy (hunter68):

Part A The first part of Ray and Kelsey's roller coaster is a curved pattern that can be represented by a polynomial function. Ray and Kelsey are working to graph a third-degree polynomial function that represents the first pattern in the coaster plan. Ray says the third-degree polynomial has 4 intercepts. Kelsey argues the function can have as many as 3 zeros only. Is there a way for the both of them to be correct? Explain your answer. No because polynomials cannot have anything that equals zero. Polynomial functions have to equal a number bigger than zero. Kelsey has a list of possible functions. Pick one of the g(x) functions below and then describe to Kelsey the key features of g(x), including the end behavior, y-intercept, and zeros. choice g(x) = x3 − x2 − 4x + 4 g(x) = x3 + 2x2 − 9x − 18 g(x) = x3 − 3x2 − 4x + 12 g(x) = x3 + 2x2 − 25x − 50 g(x) = 2x3 + 14x2 − 2x − 14 The end behavior of this g(x) function is a positive 4. The y-intercept of this function is Create a graph of the polynomial function you selected from Question 2. Part B The second part of the new coaster is a parabola. Ray needs help creating the second part of the coaster. Create a unique parabola in the pattern f(x) = ax2 + bx + c. Describe the direction of the parabola and determine the y-intercept and zeros. The safety inspector notes that Ray also needs to plan for a vertical ladder through the center of the coaster's parabolic shape for access to the coaster to perform safety repairs. Find the vertex and the equation for the axis of symmetry of the parabola, showing your work, so Ray can include it in his coaster plan. Create a graph of the polynomial function you created in Question 4. Part C Now that the curve pieces are determined, use those pieces as sections of a complete coaster. By hand or by using a drawing program, sketch a design of Ray and Kelsey's coaster that includes the shape of the g(x) and f(x) functions that you chose in the Parts A and B. You do not have to include the coordinate plane. You may arrange the functions in any order you choose, but label each section of the graph with the corresponding function for your instructor to view. Part D Create an ad campaign to promote Ray and Kelsey's roller coaster. It can be a 15-second advertisement for television or radio, an interview for a magazine or news report, or a song, poem, or slideshow presentation for a company. These are just examples; you are not limited to how you prepare your advertisement, so be creative. Make sure to include a script of what each of you will say if you are preparing an interview or a report. The purpose of this ad is to get everyone excited about the roller coaster.

OpenStudy (hunter68):

i answered Part A but i have no idea if im doing this right?

OpenStudy (hunter68):

@escamer @ddaum03

OpenStudy (hunter68):

@skatergirl147 @rarecigs

OpenStudy (hunter68):

@jovanny21 can u help?

OpenStudy (hunter68):

@drawwithapurpose

OpenStudy (hunter68):

@jr1998789

OpenStudy (hunter68):

@thalya_15

OpenStudy (hunter68):

@Trevor12

OpenStudy (drawwithapurpose):

https://www.desmos.com/calculator this might help with the rest of your questions but i don't understand part c.

OpenStudy (hunter68):

how can i get a moderater?

OpenStudy (drawwithapurpose):

a moderator? well...you just tag them like any other person

OpenStudy (drawwithapurpose):

like @mathmale

OpenStudy (hunter68):

can u help with that?

OpenStudy (drawwithapurpose):

he's a mod

OpenStudy (hunter68):

and we wait?

OpenStudy (drawwithapurpose):

mhm or ask different mods like @tkhunny

OpenStudy (hunter68):

how do i know whos a mod and whos not?

OpenStudy (drawwithapurpose):

mods are the people with purple usernames and purple boxes around their profile pictures.

OpenStudy (hunter68):

@SeanStewart16 help?

OpenStudy (drawwithapurpose):

you can also try @perl . perl is a qualified helper and a mod...so she might help

OpenStudy (hunter68):

@NetflixAndNo

OpenStudy (hunter68):

@Awolflover1

OpenStudy (hunter68):

@alibaby

OpenStudy (hunter68):

@jesus14

OpenStudy (anonymous):

yes? im so confused what do u need help in

OpenStudy (hunter68):

algebra 1

OpenStudy (hunter68):

all of this

OpenStudy (hunter68):

@PicklePlanet

OpenStudy (hunter68):

help?

OpenStudy (seanstewart16):

i know a moderator, want me to call him?

OpenStudy (hunter68):

i need help with this

OpenStudy (seanstewart16):

ill call him

OpenStudy (jesus14):

yea man this is pretty hard

OpenStudy (hunter68):

Can someone help?

OpenStudy (seanstewart16):

@mathmale

OpenStudy (hunter68):

tagged him already he hasnt come yet

OpenStudy (seanstewart16):

dude, what class is this for?

OpenStudy (hunter68):

Algebra 1

OpenStudy (anonymous):

@Awolflover1 @hartnn @hacker89 @HarleyQuinn98 @Kitten_is_back @ParthKohli @Safa102 @SamsungFanBoy @TheCuteOne25 @YoungStudier @you_me @Zeronknight

OpenStudy (seanstewart16):

@mathmale

OpenStudy (seanstewart16):

hes coming, he just noted me

OpenStudy (mathmale):

First, I need to know where you (Hunter68) currently are in your approach to this problem. Your post covers a lot of material. What specifically do you need to learn and do?

OpenStudy (hunter68):

i need help with all of it @mathmale and i am in the middle of A

OpenStudy (hunter68):

i have no idea if i am doing this right...

OpenStudy (mathmale):

"No because polynomials cannot have anything that equals zero. Polynomial functions have to equal a number bigger than zero." I can't agree to either response. First of all, we routinely set polynomials = to zero when the goal is to find roots / solutions / zeros. Secondly, polynomials can have both positive and negative values, as well as zero. I'd suggest you find a couple of third degree polynomials and graph them before you try to answer Part A.

OpenStudy (mathmale):

I've just given you some feedback. See my previous message. Are you learning from a textbook? Do you have printed study materials? If no, what is your primary source of info on Algebra I?

OpenStudy (hunter68):

I am homeschooled

OpenStudy (hunter68):

with FLVS

OpenStudy (mathmale):

So, if you were looking up "polynomials," or "graphing polynomials," where would you go? I'm glad for the opportunity to help you with this problem, but do not feel it'd be appropriate to start from scratch explaining this material. That's why I ask you to formulate specific questions that we can discuss.

OpenStudy (hunter68):

I use Desmos or Google or this

OpenStudy (mathmale):

Have you Googled "graphs of third-order polynomials?"

OpenStudy (hunter68):

I googled can 3rd order polynomials have zeros and got that

OpenStudy (mathmale):

That's entirely true, but not enough info. I looked up "third order polynomials" and obtained the following results: https://www.google.com/search?sourceid=chrome-psyapi2&ion=1&espv=2&es_th=1&ie=UTF-8&q=third-order%20polynomials&oq=third-order%20polynomials&aqs=chrome..69i57j0l5.3763j0j7 The reason I'm asking you to do some research here is that you'd benefit from seeing examples of the graphs of third-order polynomials. Can you use the Draw utility to sketch y=x^3?

OpenStudy (mathmale):



OpenStudy (hunter68):

what do u mean by sketch? and draw utility?

OpenStudy (mathmale):

I've just demonstrated it by typing / drawing y = x ^3 and by setting up a system of coordinate axes (x- and y-axes).

OpenStudy (mathmale):

Please have a look at http://www.sscc.edu/home/jdavidso/Math/Catalog/Polynomials/Third/Third.html because this site features graphs of a couple of third-order polynomials. It's essential that y ou know what such polynomials look like.

OpenStudy (mathmale):

You could edit my graph by clicking on the pencil icon in the upper, right hand corner of the graph. A new graph in a different color will open. You can type or draw on that graph. Graph y=x^3.

OpenStudy (hunter68):

im more confused than ever

OpenStudy (mathmale):

Please look up this web page. Do you see examples of graphs of 3rd order polynomials there?

OpenStudy (hunter68):

nvm im going to see if Ms.Roessler can help...(my teacher) i am sorry for the inconvenience but your just confusing me more

OpenStudy (mathmale):

All right. Sorry that our interaction is not working out.

OpenStudy (hunter68):

its ok openstudy is awesome :)

OpenStudy (mathmale):

It may sound old-fashioned, but asking Ms. Hunter about a good algebra book that you could purchase and study might be one of the best things you could do to help yourself. Back in my day (1944) there was no such thing as an "Internet" or "OpenStudy," and so everything I learned came from books. Just a thought. I get the impression that you need more background in algebra than you currently have, to do the problem you've shared here on OpenStudy. Good luck, Hunter. Let me know what you learn from Ms. Hunter.

OpenStudy (hunter68):

Ms.Roessler* and please call me Ray and thanks

OpenStudy (mathmale):

OK. Hope Ms. R. can help you; sometimes it's best to have a live teacher who gets you involved. Take care, Ray. Bye. Over and out.

OpenStudy (hunter68):

bye

OpenStudy (hunter68):

ok now i need help with part C and D

OpenStudy (hunter68):

@SeanStewart16

OpenStudy (hunter68):

@ilovebmth1234